OpenStudy (anonymous):

Show algebraically and graphically that the Squaring Function and the Square Root Function are inverses of each other.

OpenStudy (anonymous):

unfortunately they aren't

OpenStudy (anonymous):

if \(x=-3\) then \(x^2=(-3)^2=9\) and \(\sqrt{9}=3\)

OpenStudy (anonymous):

they are only inverse if you restrict the domain of \(f(x)=x^2\) to \([0,\infty)\)

OpenStudy (anonymous):

ok so how would you algebraically show that when the domain is restricted

OpenStudy (anonymous):

if \(x\geq\) then \((\sqrt{x})^2=x\) and \(\sqrt{x^2}=x\)

OpenStudy (anonymous):

thanks